### v1

# Hubble tension tomography: BAO vs SnIa distance tension

Abstract
We investigate the redshift dependence of the Hubble tension by comparing the luminosity distances obtained using an up-to-date BAO dataset (including the latest DESI data) calibrated with the CMB-inferred sound horizon, and the Pantheon+ SnIa distances calibrated with Cepheids. Using a redshift tomography method, we find: 1) The BAO-inferred distances are discrepant with the Pantheon+ SnIa distances across all redshift bins considered, with the discrepancy level varying with redshift. 2) The distance discrepancy is more pronounced at lower redshifts ($z\in[0.1,0.8]$) compared to higher redshifts ($z\in[0.8,2.3]$). The consistency of $\Lambda$CDM best fit parameters obtained in high and low redshift bins of both BAO and SnIa samples is investigated and we confirm that the tension reduces at high redshifts. Also a mild tension between the redshift bins is identified at higher redshifts for both the BAO and Pantheon+ data with respect to the best fit value of $H_{0}$ in agreement with previous studies which find hints for an ’evolution’ of $H_{0}$ in the context of $\Lambda$CDM . These results confirm that the low redshift BAO and SnIa distances can only become consistent through a re-evaluation of the distance calibration methods. An $H(z)$ expansion rate deformation alone is insufficient to resolve the tension. Our findings also hint at a possible deviation of the expansion rate from the Planck18/$\Lambda$CDM model at high redshifts $z\gtrsim 2$. We show that such a deformation is well described by a high redshift transition of $H(z)$ like the one expressed by $\Lambda_{\rm s}$ even though this alone cannot fully resolve the Hubble tension due to its tension with intermediate/low $z$ BAO data.
I Introduction
The sound horizon scale $r_{s}$ at recombination (standard ruler) and the standardized bolometric absolute magnitude $M_{B}$ of Type Ia supernovae (SnIa standard candles) have been used as probes for the measurement of cosmological distances and thus for the measurement of the Hubble constant $H_{0}$, the most fundamental parameter of cosmology. The best fit values of $H_{0}$ obtained using the two distance calibrators are at $5\sigma$ discrepancy with each other. This is the most important challengePerivolaropoulos and Skara (2022a); Abdalla et al. (2022); Di Valentino et al. (2021) for the current standard cosmological model the $\Lambda CDM$.
Each type of measurement makes specific assumptions whose possible violation would lead to significant systematic errors that would invalidate the accuracy of the corresponding measurement. Measurements based on SnIa standard candles assume the validity of the distance ladder approach and in particular that physical laws and environmental effects around calibrated SnIa are the same in all the three rungs of the distance ladder. Measurements based on the sound horizon scale used as a standard ruler assume that the physical laws before recombination are consistent with the standard cosmological model and that the expansion history of the universe $H(z)$ is consistent with the standard $\Lambda$CDM model and its parameters determined by the Planck18Aghanim et al. (2020) (Planck18/$\Lambda$CDM ).
In accordance with the above three types of assumptions there are three classes of models for the resolution of the Hubble tension:
  • Ultralate time models: These modelsMarra and Perivolaropoulos (2021); Alestas et al. (2021a); Perivolaropoulos and Skara (2021) assume that there is a physics or environmental change between the second and third rungs of the distance ladder ie that the Cepheid calibrated SnIa are not the same as the Hubble flow SnIa either due to a change of the physical laws or due to some change in their structure or environment (dust, metalicity etc). Thus, $M_{B}$ in the third rung of the distance ladder (Hubble flow rung $z\in[0.01,0.1]$) is lower than the value measured in the second rang. Thus these models explore the degeneracy between $H_{0}$ and $M_{B}$ in the context of the Hubble flow observable
    ${\cal M}\equiv M_{B}+5log\frac{c/H_{0}}{Mpc}+25$ (1)
    to lead to a lower $H_{0}$ measurement in the context of the distance ladder measurements.
    The main problem of this class of models is fine tuning. There is no current clear theoretical motivation for the assumed physics transition at such ultralate redshifts ($z\sim 0.01$) even though it is straightforward to construct such models using a degree of fine tuningPerivolaropoulos and Skara (2022b). On the other hand, the physics transitions hypothesis is easily testable using a wide redshift data and in fact there are some hints for such an affect in both Cepheid and Tully-Fisser data Perivolaropoulos and Skara (2022c); Alestas et al. (2021b).
  • Early time models: These models assume physics beyond the standard model to decrease the sound horizon scale at recombination
    $r_{s}=\int_{z_{rec}}^{\infty}\frac{dz\;c_{s}(z)}{H(z;\Omega_{0b}h^{2},\Omega_{0\gamma}h^{2},\Omega_{0CDM}h^{2})}$ (2)
    to a lower value induced by Early Dark Energy Poulin et al. (2019); Kamionkowski and Riess (2023); Simon et al. (2023); Braglia et al. (2020); Niedermann and Sloth (2020); Smith et al. (2021) (EDE), modified gravityBraglia et al. (2021); Brax et al. (2014); Adi and Kovetz (2021); Clifton et al. (2012); Lin et al. (2019); Di Valentino et al. (2016); Akrami et al. (2021) or radiationSeto and Toda (2021); Sakstein and Trodden (2020). Thus, these models exploit the degeneracy of $r_{s}$ with $H_{0}$ in the context of the observable angular scale of the CMB sound horizon $r_{s}$
    $\theta_{s}=\frac{r_{s}H_{0}}{\int_{0}^{z_{rec}}1/E(z)}$ (3)
    to lead to a higher value of $H_{0}$ while respecting the best fit Planck18/$\Lambda$CDM form of $E(z)\equiv H(z)/H_{0}$ between recombination and present time.
    The main problem of this class is that despite fine tuning of theoretical models that support them, they are only able to decrease the statistical significance of the Hubble tension and they can not fully eliminate it. In addition, they favor a higher value of matter density parameter $\Omega_{0m}$ thus worsening the $S_{8}$ tension Alestas and Perivolaropoulos (2021); Jedamzik et al. (2021); Vagnozzi (2023).
  • Late time models ($H(z)$ deformation): These models assume that there is deformation of the Hubble expansion history $H(z)$ with respect to the Planck18/$\Lambda$CDM prediction at late cosmological times ($z\lesssim 2$). This deformation shifts the value of $H_{0}$ measured by the sound horizon scale standard ruler to a value consistent with the distance ladder measurement by effectively changing the denominator of eq. (3). However, as discussed in the present analysis even this model has difficulty to fit simultaneously BAO and SnIa data.
    The main problem of this class of models is that $E(z)\equiv H(z)/H_{0}$ deformations are highly constrained by BAO and SnIa data in redshifts larger than $z\simeq 0.1$ Alestas et al. (2022); Alestas and Perivolaropoulos (2021); Alestas et al. (2020); Brieden et al. (2022); Keeley and Shafieloo (2023); Clark et al. (2021); Chen et al. (2021); Anchordoqui et al. (2022); Mau et al. (2022); Cai et al. (2022); Heisenberg et al. (2022); Vagnozzi et al. (2022); Davari and Khosravi (2022). Thus, this class of models have been put in disfavor during the last few years. Exception constitutes the $\Lambda_{\rm s}$ model which shifts the deformation to high redhifts ($z\simeq 2$) and makes it abrupt so that it has a minimal effect on late data while maintaining consistency with the distance to the CMB sound horizon at last scattering Akarsu et al. (2023a).
The $H(z)$ deformation models are designed to be consistent with three types of cosmological data.
  • Low $z$ ($z\lesssim 0.1$) calibrated Hubble flow data (eg the SH0ES data Riess et al. (2022)) which measure $H_{0}=73.04\pm 1km/secMpc^{-1}$.
  • The distance to the CMB sound horizon scale $r_{s}$ as measured by the Planck CMB power spectrum.
  • The distance to BAO and SnIa data as obtained with the corresponding calibrators (CMB $r_{s}$ for BAO and Cepheid $M_{B}$ for SnIa).
These models perform very well in fulfilling the first two requirements. However, they all have problems fulfilling the third requirement. An interesting questing thus emerges
Is this difficulty due to the fact that there is a specific feature missed by the $H(z)$ parametrizations considered so far? Or there is a generic inconsistency between the BAO and SnIa data when the proper calibration is used in each case?.
This question is addressed in the present analysis using a redshift tomography method. We thus test the consistency of the cosmological distances obtained with the properly calibrated BAO and SnIa data with each other and with the Planck18/$\Lambda$CDM $E(z)$ in various redshift bins. We thus identify those redshifts where the distances obtained with the two classes of data are consistent and search for redshifts where there is a generic inconsistency between these distances. For the later redshift range it would be impossible to fit simultaneously the BAO and SnIa for any parametrization $H(z)$ due to the generic inconsistency between the two classes of data.
We thus find the level of inconsistency between $r_{s}$ calibrated BAO distances and Cepheid $M_{B}$ calibrated SnIa measured distances as a function of redshift. This may be interpreted as the level of the Hubble tension as a function of redshift. We also find the level of consistency of the measured distances with the predictions of Planck18/$\Lambda$CDM as a function of redshift.
Recent studies have investigated the redshift dependence of the Hubble tension using various probes and methods Dainotti et al. (2021); Colgáin et al. (2022); Krishnan et al. (2021). These analyses consistently find a reduction in the tension at higher redshifts, hinting at a possible low-redshift origin for the discrepancy. However, the specific redshift range and significance of this effect varies between studies, motivating further investigation.
Previous studies Pogosian et al. (2022) have demonstrated a general trend for inconsistency between the cosmological distances obtained with BAO and SnIa data. We quantify this finding using redshift tomography and a more extensive up-to-date BAO dataset. Promising approaches like $\Lambda_{\rm s}$ Akarsu et al. (2020, 2023a) are also investigated, which introduce a rapid transition in $H(z)$ at high redshifts to improve consistency with SnIa data. However, we find that the fundamental tension between BAO and SnIa data creates problems even for these models.
Potential systematic effects that could influence the redshift dependence of the Hubble tension are an important consideration. For example, selection effects in the SnIa samples at different redshifts Brout et al. (2019) or evolution in the SnIa luminosity with redshift Tutusaus et al. (2019) could impact the tension measurement. On the BAO side, the determination of the sound horizon scale $r_{d}$ relies on assumptions about early universe physics that could potentially be violated Jedamzik and Pogosian (2020). While a full investigation of these effects is beyond the scope of this paper, they are important caveats to keep in mind when interpreting the results.
The structure of this paper is the following: In section II we present the method used in our analysis which involves binning distance measurements using BAO data and comparing them with the corresponding distances obtained with binned Pantheon+ SnIa data. This tomography is used for estimating the statistical level of the calibrator mismatch as a function of redshift (Hubble tension tomography). Our results are presented in section III. Finally, section IV we conclude, summarize and discuss possible extensions of our analysis.
II Method: Binning distance measurements with Pantheon+ and BAO
The Pantheon+ sample provides SnIa luminosity distance $d_{L}$ and distance moduli $\mu$ measurements for redshifts in the range $z\in[0.001,2.3]$. These measurements are calibrated by the second rung of the distance ladder using Cepheids measuring the SnIa absolute magnitude as $M_{B}=-19.25\pm 0.01$ Riess et al. (2022); Perivolaropoulos and Skara (2022c, 2023).
Assuming validity of the distance duality relation $d_{L}=(1+z)^{2}D_{A}$, the BAO data calibrated by the comoving CMB sound horizon scale standard ruler $r_{d}=147.1Mpc$ at the end of the baryonic drag epoch Aghanim et al. (2020), also lead to measurements of the luminosity distance $d_{L}$ and distance moduli $\mu$, in the redshift range $z\in[0.1,2.4]$. The CMB-inferred sound horizon scale $r_{d}$ is determined by fitting the standard $\Lambda$CDM model to the Planck CMB power spectra data, assuming standard pre-recombination physics Aghanim et al. (2020). It corresponds to the comoving size of the sound horizon at the end of the baryonic drag epoch, when photon pressure no longer supports acoustic oscillations in the primordial plasma (see Aubourg et al. (2015) for a review). The consistency level of these measurements with the corresponding measurements of the Pantheon+ sample as a function of redshift bins reflects the level of the Hubble tension in redshift space. The method to perform this Hubble tension tomography is described in this section.
II.1 Binning the distance moduli residual of Pantheon+
The Pantheon+ sampleBrout et al. (2022a); Scolnic et al. (2022) consists of a table of 1701 rows (plus a header) and 47 columns. A $1701\times 1701$ covariance matrix is part of it and it represents the covariance between all the SnIa due to systematic and statistical uncertainties of the distance moduli. In this analysis the information we need to utilize from the sample is:
  • The Hubble diagram redshift (with CMB and peculiar velocity corrections)
  • The corrected apparent magnitude of the SnIa as well as its uncertainty which includes errors due to peculiar velocity
  • The distance moduli $\mu$ calculated by subtracting the absolute magnitude of the SnIa $M_{SH0ES}=-19.253$ from the apparent magnitude and also their uncertainties obtained through the diagonal of the covariance matrix
  • The distance moduli $\mu_{Ceph}$, as found by the SH0ES distance ladder analysis, of the SnIa in Cepheid hosts with their uncertainties being incorporated in the covariance matrix
The data can be found in the github repository in this url.
The SnIa distance moduli of the Pantheon+ sample measured in the Hubble flow are used to constrain the luminosity distance $d_{L}$ through the relation
$\mu(z)=m(z)-M_{B}=5log(d_{L}(z)/Mpc)+25$ (4)
where the luminosity distance is connected with the Hubble expansion rate:
$d_{L}(z)=(1+z)c\int^{z}_{0}\frac{dz^{\prime}}{H_{0}\;E(z^{\prime})}$ (5)
$c$ is the speed of light and the dimensionless Hubble parameter $E$ in the context of $\Lambda$CDM is given by:
$E=\sqrt{(a^{-3}\Omega_{0m}\left(1+\frac{a_{eq}}{a}\right)+\left(1-\Omega_{0m}(1+a_{eq})\right)}$ (6)
with $a_{eq}$ the scale factor corresponding to the time when matter and radiation were equal
$a_{eq}=\frac{1}{(z_{eq}+1)},z_{eq}=2.5\cdot 10^{4}\Omega_{0m}h^{2}(T_{CMB}/2.7)^{-4}$ (7)
with $T_{CMB}=2.7255K$.
To get the theoretical distance modulus corresponding to the Planck18/$\Lambda$CDM model we simply need to incorporate the Planck 2018 values for the present matter density parameter $\Omega_{0m}=0.3166$ and the Hubble constant $H_{0}=67.4\quad km\quad s^{-1}\quad Mpc^{-1}$ ($h=0.674$) into equations (4)-(6) to obtain $\mu_{P\Lambda CDM}(z)$.
We define the distance modulus residual between the theoretical values of the Planck18/$\Lambda$CDM model and the observational data obtained through the Pantheon+ sample:
$\Delta\mu(z_{i})\equiv\mu_{obs}(z_{i})-\mu_{P\Lambda CDM}(z_{i})$ (8)
We have binned this residual in 13 evenly spaced bins for $z\in[0.001,2.3]$ to compare with corresponding BAO data. The redshift for each bin is the mean of all redshifts within said bin. For the calculation of the distance modulus for each bin the covariance matrix was incorporated. This was done by first constructing a vector for each bin $j$ with components
$Q_{i}=\Delta\mu(z_{i})-\mu_{cj}$ (9)
Components of $Q$ out of the $j$ bin were set to 0. In (9), $\mu_{cj}$ is the residual distance modulus in the $j^{th}$ bin and $\Delta\mu(z_{i})$ is defined by eq. (8) for a given redshift $z_{i}$ in the redshift range of the corresponding bin. Then to obtain the best fit value of $\Delta\mu(z_{Meanj})$ (with $z_{Meanj}$ being the mean redshift in the redshift range of the $j^{th}$ bin), we found the value of the best fit $j^{th}$ bin distance modulus residual $\mu_{cj}$ that minimizes the appropriate $\chi^{2}$ function:
$\chi^{2}_{j}=Q^{T}\cdot C^{-1}\cdot Q$ (10)
where $C^{-1}$ is the inverse of the covariance matrix of the Pantheon+ sample. Using the definition (9) in (10) the proper part of the inverse covariance matrix corresponding to the $j^{th}$ bin is automatically selected. In this fashion we end up with 13 different Planck18/$\Lambda$CDM residual distance moduli values $\mu_{cj}$, one for each bin. The uncertainties for each bin $j$ were calculated in the standard way using the inverse of the Fisher matrix
$F=\frac{1}{2}\frac{\partial^{2}\chi^{2}(\mu_{cj})}{\partial\mu_{cj}^{2}}$ (11)
We have also split the Pantheon+ sample into two larger bins for redshifts $0.80.1<z<$ 0.8 and $2.30.8<z<$ 2.3 in order to fit for each bin the $\Lambda$CDM cosmological parameters and identify the self consistency of the Pantheon+ sample and also compare with the corresponding best fit parameters of a BAO sample discussed in the next subsection. For each bin a data vector can be constructed as follows:
$Q^{\prime}_{i}=\begin{cases}m_{B,i}-M_{B}-\mu^{Ceph}_{i}&,\text{Cepheid hosts}\\ m_{B,i}-M_{B}-\mu_{\Lambda CDM}(z_{i})&,\text{Otherwise}\\ \end{cases}$ (12)
where in each bin we have also included the Cepheid host SnIa distance moduli as found by the SH0ES distance ladder analysis. For each bin we evaluate the $\chi^{2}$ function:
$\chi^{\prime 2}_{bin}=Q^{\prime T}\cdot C^{-1}\cdot Q^{\prime}$ (13)
and minimize it to obtain the best fit values for $M_{B}$, $\Omega_{0m}$ and $h$ and the corresponding likelihood contours discussed in the next section. We have also constructed the distance moduli residual for the two new large bins as described above for the smaller bins. These distance moduli residuals and $\Lambda$CDM parameter contours were then compared with the corresponding BAO quantities described in the next subsection.
II.2 Distance moduli from BAO
We have compiled a list of transverse BAO measurements, of the rescaled comoving angular diameter distance $\frac{D_{M}}{r_{d}}$, where $r_{d}$ is the sound horizon scale during the drag epoch in which the photon drag on regular matter (baryons) becomes negligible. The sample created includes both data as recent as the DESI 2024 Adame et al. (2024) as well as some earlier measurements including the latest releases of the Sloan Digital Sky Survey (SDSS) and the Dark Energy Survey (DES). The full sample we use is shown in Table 1
Table 1: $D_{M}/r_{d}$ measurements and the corresponding references. The sample consists of 33 measurements. Measurements with an asterisk are the $D_{M}/r_{d}$ value calculated using the angular diameter distance $D_{A}=D_{M}/(1+z)$ and the fiducial $r_{d}^{fid}$ value $r_{d}=147.18Mpc$ obtained by Planck18 in the context of $\Lambda$CDM . More details about the construction of the * datapoints are presented in the Appendix. V.
N $z_{eff}$ $D_{m}/r_{d}$ References Year
1 0.32* 8.54±0.25 Chuang et al Chuang et al. (2017) 2016
2 0.32 8.76±0.14 BOSS collaboration Alam et al. (2017) 2016
3 0.38 10.27±0.15 BOSS collaboration Alam et al. (2017) 2016
4 0.44* 11.79±1.11 Blake et al Blake (2012) 2012
5 0.51 13.38±0.18 BOSS collaboration Alam et al. (2017) 2016
6 0.51 13.62±0.25 DESI Collaboration Adame et al. (2024) 2024
7 0.54* 14.76±0.68 Seo et al Seo (2012) 2012
8 0.57 14.74±0.24 BOSS collaboration Alam et al. (2017) 2016
9 0.59* 15.29±0.25 Chuang et al Chuang et al. (2017) 2016
10 0.60* 14.99±1.04 Blake et al Blake (2012) 2012
11 0.61 15.45±0.22 BOSS collaboration Alam et al. (2017) 2016
12 0.70* 17.28±0.89 Sridhar et al Sridhar et al. (2020) 2020
13 0.70 17.96±0.51 Zhao Gong-bo et al Zhao et al. (2021) 2021
14 0.71 16.85±0.32 DESI Collaboration Adame et al. (2024) 2024
15 0.73* 18,03±1.26 Blake et al Blake (2012) 2012
16 0.77 18.85±0.38 Wang et al Wang et al. (2020) 2020
17 0.80 19.54±2.07 Zhu et al Zhu et al. (2018) 2018
18 0.84 19.15±0.58 Chan et al Chan et al. (2022) 2022
19 0.85 19.5±1.0 Tamone et alTamone et al. (2020), de Mattia et alde Mattia et al. (2021) 2020
20 0.87* 21.39±1.39 Sridhar et al Sridhar et al. (2020) 2020
21 0.93 21.71±0.28 DESI Collaboration Adame et al. (2024) 2024
22 1.00 23.13±2.07 Zhu et al Zhu et al. (2018) 2018
23 1.32 27.79±0.69 DESI Collaboration Adame et al. (2024) 2024
24 1.48 30.21±0.79 Hou et alHou et al. (2020), Neveux et alNeveux et al. (2020) 2020
25 1.50 30.51±1.85 Zhu et al Zhu et al. (2018) 2018
26 2.00 36.18±1.69 Zhu et al Zhu et al. (2018) 2018
27 2.20 38.09±1.73 Zhu et al Zhu et al. (2018) 2018
28 2.33 37.5±1.1 Du Mas des Bourbuox et aldu Mas des Bourboux et al. (2020) 2020
29 2.33 39.71±0.94 DESI Collaboration Adame et al. (2024) 2024
30 2.34* 37.72±2.18 Delubac et al Delubac et al. (2015) 2014
31 2.35 36.3±1.8 Blomqvist et al Blomqvist et al. (2019) 2019
32 2.40 36.6±1.2 De Mas des Bourboux et al du Mas des Bourboux et al. (2017) 2017
The distance modulus residual can now be written as:
$\begin{split}\Delta\mu(z_{i})&=\mu_{obs}(z_{i})-\mu_{P\Lambda CDM}(z_{i})\\ &=5log\left(\frac{d_{L}^{obs}(z_{i})}{d_{L}^{P\Lambda CDM}(z_{i})}\right)\end{split}$ (14)
The ratio of the observed luminosity distance over the Planck 2018 theoretical luminosity distance, $\frac{d_{L}^{obs}(z_{i})}{d_{L}^{P\Lambda CDM}(z_{i})}$, can be replaced by:
$\frac{d_{L}^{obs}(z_{i})}{d_{L}^{P\Lambda CDM}(z_{i})}=\frac{D_{M}^{obs}(z_{i})}{D_{M}^{P\Lambda CDM}(z_{i})}$ (15)
since the comoving angular distance $D_{M}$ is simply $D_{M}=(1+z)D_{A}$, where $D_{A}$ is the angular diameter distance which in turn is $D_{A}=d_{L}/(z+1)^{2}$. Therefore, the distance moduli residuals can be written as:
$\Delta\mu(z_{i})=5log\left(\frac{D_{M}^{obs}(z_{i})}{D_{M}^{P\Lambda CDM}(z_{i})}\right)$ (16)
where we have estimated $D_{M}^{obs}(z_{i})$ by multiplying the entries of Table 1 by the best fit value $r_{d}=147.18Mpc$ from Planck18Aghanim et al. (2020). For the theoretical Planck18/$\Lambda$CDM $D_{M}(z_{i})$ we have incorporated the Planck 2018 $\Lambda$CDM values for the Hubble constant and the present matter density parameter as mentioned above in section II.1. We also assume flat universe and thus use the present $\Lambda$ density parameter $\Omega_{0\Lambda}=0.685$, the present radiation density parameter $\Omega_{0r}=9.26*10^{-5}$ and the sound horizon drag scale $r_{d}=147.18$:
$D_{M}^{P\Lambda CDM}(z_{i})=\\ \int^{z_{i}}_{0}\frac{c}{H_{0}\sqrt{\Omega_{0m}(1+z)^{3}+\Omega_{0\Lambda}+\Omega_{0r}(1+z)^{4}}}$ (17)
The distance modulus residual $\Delta\mu(z_{i})$ can then be plotted as a function of redshift $z_{i}$ (left panel of Fig. 1). The uncertainties are calculated using the error propagation method as follows:
$\sigma_{i}=\frac{5}{ln10}\frac{1}{D_{M}^{obs}(z_{i})}\cdot\sigma_{D_{M}^{obs}}(z_{i})$ (18)
where we have omitted the effect of the uncertainty of $r_{d}$ because it is easy to show that it is subdominant. We have also merged some of the points which have the same effective redshift value. This was done by first calculating the weighted $D_{M}$ ratios Press et al. (2007):
$D_{M,weighted}=\frac{\sum D_{Mi}/\sigma_{i}}{\sum 1/\sigma_{i}}$ (19)
and then calculating the combined uncertainty:
$\sigma_{combined}=\frac{1}{\sqrt{\sum 1/\sigma_{i}}}$ (20)
For the construction of the right panel of Fig. 1, we have also considered three redshift bins for redshift ranges $0.80.1<z<$ 0.8, $2.00.8<z<$ 2.0 and $2.52.0<z<$ 2.5. For each bin we get the distance modulus residuals $\Delta\mu_{i}$ as shown in Eq. (14) and the uncertainties $\sigma_{i}$ for each residual (see Eq. 18). The redshift for each bin is the mean of all the redshifts of said bin, the distance modulus residual is the weighted mean:
$\Delta\mu_{bin}=\frac{\sum\frac{\Delta\mu_{i}}{\sigma^{2}_{i}}}{\sum 1/\sigma_{i}^{2}}$ (21)
and the uncertainty is the combined uncertainty:
$\sigma_{bin}=\sqrt{\frac{1}{\sum 1/\sigma_{i}^{2}}}$ (22)
The bin residulal distance moduli are shown in the right panel of Fig. 1 along with the two Pantheon+ bins discussed at the end of section II.1 for a comparison to be made. The comparison is discussed in the next section.
For the comparison of the best fit cosmological parameters $h$ and $\Omega_{0m}$ with Pantheon+, we have constructed two redshift bins in the same redshift ranges as those used for the Pantheon+ data ($0.80.1<z<$ 0.8 and $2.50.8<z<$ 2.5  [^1] It should be noted here that the redshift range of the second bin goes beyond redshift 2.3, up to redshift 2.5 since we have BAO data for these redshifts as well..
Thus, for each bin we construct the $\chi^{2}$ function as follows:
$\chi^{2}=\sum\frac{(D_{M}^{obs}(z_{i})-D_{M}^{\Lambda CDM}(z_{i}))^{2}}{\sigma_{D_{M}^{obs}}^{2}(z_{i})+\sigma_{r_{d}}^{2}}$ (23)
where we estimate $(D_{M}^{obs}(z_{i})$ using the Planck18 best fit value for $r_{d}=147.18Mpc$. By minimizing this function we obtain the best fit values for $h$ and $\Omega_{0m}$. The parameter uncertainties were again calculated using the Fisher matrix:
$F_{ij}=\frac{1}{2}\frac{\partial^{2}\chi^{2}(p)}{\partial p_{i}\partial p_{j}}$ (24)
with the two parameters $p_{i}$ and $p_{j}$ being $\Omega_{0m}$ and $h$. The uncertainties in this case are the square roots of the diagonal elements of the inverse Fisher matrix. The corresponding contour plots can then be constructed to represent these results.
II.3 Model Residuals for the $wCDM$ and $\Lambda_{\rm s}$ Models
The study of $H(z)$ deformation models is driven by the need for a cosmological framework that is in agreement with Cosmic Microwave Background (CMB) observations at high redshifts ($z\simeq 1100$) as well as with local measurements of the Hubble constant ($H_{0}$). These models are designed to maintain the angular diameter distance to the sound horizon scale at recombination as observed by the Planck18/$\Lambda$CDM model, thereby ensuring consistency with the CMB power spectrum. They also assume that $\Omega_{0m}h^{2}$ has the same value as the one measured by the CMB angular power spectrum while setting constraining $h$ by local measurements Alestas et al. (2020). Interesting such models are the $wCDM$ Alestas et al. (2020) and $\Lambda_{\rm s}$ Akarsu et al. (2020, 2023a) models investigated for their $H(z)$ deformation approach to the $H_{0}$ tension.
wCDM Model: The $wCDM$ model modifies the standard equation of state for dark energy, $w$, from the Planck18/$\Lambda$CDM value of $-1$. The Hubble parameter for this model is given by
$\left(\frac{H}{H_{0}}\right)^{2}=\Omega_{r0}(1+z)^{4}+\Omega_{m0}(1+z)^{3}+\Omega_{DE0}(1+z)^{3(1+w)}.$ (25)
$\boldsymbol{\Lambda sCDM}$ Model: Alternatively, the $\Lambda_{\rm s}$ model posits a step-like alteration in the dark energy density at a transition redshift $z_{*}$. The corresponding Hubble parameter is modeled as
$\left(\frac{H}{H_{0}}\right)^{2}=\Omega_{r0}(1+z)^{4}+\Omega_{m0}(1+z)^{3}+\Omega_{\Lambda s0}\text{sgn}(z-z_{*}),$ (26)
where sgn represents the signum function, indicative of the dark energy density change.
In this context, the residual distance modulus, $\Delta\mu(z)$, quantifies the difference between the distance modulus as predicted by a given cosmological model, $\mu_{\text{model}}(z)$, and the distance modulus according to the Planck18 $\Lambda$CDM standard model, $\mu_{\text{Planck18/$\Lambda$CDM }}(z)$. It is defined as:
$\Delta\mu(z)=\mu_{\text{model}}(z)-\mu_{\text{Planck18/$\Lambda$CDM }}(z).$ (27)
where ’model’ may correspond eg to $wCDM$ or to $\Lambda sCDM$. This residual is a critical diagnostic tool to assess the viability of models such as $wCDM$ and $\Lambda_{\rm s}$ in resolving the $H_{0}$ tension, by comparing their predictions with the well-established $\Lambda$CDM model as fitted to Planck data and with the intermediate SnIa and BAO data.
For the $wCDM$ model, with $w\simeq-1.2$, there is a successful interpolation between the Cepheid-calibrated low-$z$ SnIa measurements and the CMB-inferred distances to the sound horizon at high-$z$ Alestas et al. (2020). This is illustrated in Figure 1, where the residual distance moduli for the $wCDM$ model are compared to those of the $\Lambda_{\rm s}$ model, with the BAO and SnIa properly calibrated distance moduli residuals and with the SH0ES-Patheon+ $\Lambda$CDM Brout et al. (2022b) and Planck18/$\Lambda$CDM Aghanim et al. (2020) residual distance moduli (the later has central value at 0 by definition). The best fit $\Lambda_{\rm s}$ modelAkarsu et al. (2020, 2023a, 2023b) with parameter values obtained from Akarsu et al. (2023b) (Table I last column) ($z_{*}=1.72$), provides a better fit to the more constraining SnIa data at the expense of a worse fit to the BAO data which however have larger uncertainties and thus it leads to a better $\chi^{2}$ than $\Lambda$CDM and other $H(z)$ deformation models like $wCDM$ as shown in Fig. 1. Thus this feature or high $z$ ’break’ of $H(z)$ provides an advantage for $\Lambda_{\rm s}$ but does not overcome the fundamental problem of the $H(z)$ deformation models which is the generic inconsistency between SnIa and BAO data.
Parameter values of $H(z)$ deformation models: For the $wCDM$ model residual we use the cosmic parameters as calculated by the method described in Alestas et al. (2020). For the $w$ parameter we use the following formula:
$h(w)\simeq-0.3093w+0.3647$ (28)
where for $h$ we set $0.73$ to find $w_{model}=-1.188$. Then, for the present matter density parameter we use the Planck18/$\Lambda$CDM value
$\overline{\omega_{0m}}=0.1430\pm 0.0011$ (29)
where $\overline{\omega_{0m}}=\Omega_{0m}h^{2}$ which leads to $\Omega_{0m}=0.267$ for $H(z)$ deformation models consistent with local measurements of $h$ (eg $wCDM$).
Figure 1: The binned BAO and Pantheon+ luminosity distance moduli (residuals with respect to the Planck18/$\Lambda$CDM ). The residual distance moduli for the best fit $wCDM$ and $\Lambda_{\rm s}$ are also shown. Larger redshift bins are shown in the right panel. The inconsistency between BAO and SnIa measured distance moduli is evident. The full Pantheon+ distance moduli are also shown in the grey background. The green points correspond to the latest DESI BAO data.
Figure 2: The consistency between BAO and Pantheon+ data for the low (left) and high (right) redshift bins. Notice that the discrepancy is significantly larger for the low redshift bin.
Figure 3: The consistency between low and high redshift bins for BAO (left) and Pantheon+ (right) data. Notice that there is no evident inconsistency between the redshift bins in the context of each individual dataset.
Figure 4: The likelihood contours of $\Lambda$CDM parameters obtained with the full BAO and Pantheon+ samples. The tension between the total BAO and Pantheon+ data sets in intermediate redshifts is apparent.
Table 2: Best-fit cosmological parameters for two redshift bins within the $\Lambda$CDM model.
Redshift Bin $\boldsymbol{\Omega_{0m}}$ $\boldsymbol{h}$ $\boldsymbol{M_{B}}$
First BAO Bin ($0.80.1<z<$ 0.8) $0.3\pm 0.1$ $0.68\pm 0.03$
Second BAO Bin ($2.30.8<z<$ 2.3) $0.35\pm 0.06$ $0.67\pm 0.03$
First Pan+ Bin ($0.80.1<z<$ 0.8) $0.37^{+0.04}_{-0.06}$ $0.742^{+0.008}_{-0.01}$ $-19.21^{+0.02}_{-0.03}$
Second Pan+ Bin ($2.30.8<z<$ 2.3) $0.38^{+1}_{-0.2}$ $0.71^{+0.1}_{-0.2}$ $-19.20^{+0.7}_{-0.4}$
Total BAO $0.34\pm 0.03$ $0.67\pm 0.01$
Total Pantheon+ $0.33\pm 0.02$ $0.73\pm 0.01$ $-19.25\pm 0.03$
The luminosity distance modulus for the model is as defined in equation (4) and the luminosity distance is as defined in equation (5), with the adjustment of using $H$ as defined in equation (25).
Similarly, we construct the distance modulus residual for the $\Lambda_{\rm s}$ model Akarsu et al. (2021) with best fit parameter values as obtained in Akarsu et al. (2023b). The best fit parameter values we use for the $\Lambda_{\rm s}$ model are taken as $z_{*}=1.72$, $\Omega_{m,\Lambda sCDM}=0.2646$ and $h_{0,\Lambda sCDM}=0.731$. These are taken from Akarsu et al. (2023b), specifically from Table 1 last column.
III Results: The Hubble Tension in Redshift Bins
The focus of our analysis is the level of mismatch of the distance moduli measured using BAO with the corresponding distance moduli measured with SnIa at various redshift bins. This is mainly demonstrated in Fig. 1 which shows the distance moduli residuals with respect to Planck18/$\Lambda$CDM of the following:
  • The BAO data of the Table 1 compilation (blue points). The latest DESI data are the green points shown separately for comparison with earlier points. The full BAO data are binned in wide redshift bins in the right panel.
  • The binned Pantheon+ data (red points). These are binned in wide redshift bins in the right panel. The full Pantheon+ distance moduli are also shown in the grey background. The bin corresponding to $z<0.01$ has a much higher residual distance modulus compared to other bins due to the volumetric redshift scatter bias Perivolaropoulos and Skara (2023).
  • The Planck18/$\Lambda$CDM model with uncertainties (upper band around zero).
  • The Pantheon+ best fit $\Lambda$CDM model (lower band around -0.2).
  • The $wCDM$ model designed to interpolate between the SH0ES $H_{0}$ and the Planck18 angular diameter distance to recombination consistent with the CMB angular power spectrum data (green line)
  • The best fit $\Lambda_{\rm s}$ model with respect to Pantheon+, BAO, SH0ES and CMB data with best fit discussed in the previous section.
The following comments can be made on the results shown in Fig. 1
  • There is a clear overall tension between the distance moduli measured by BAO and SnIa data. The distance moduli Planck18/$\Lambda$CDM residuals measured by BAO appear to be systematically higher than the distance moduli residuals measured by SnIa calibrated by Cepheids but lower compared to Planck18/$\Lambda$CDM expecially at $z>1$. Similarly, the model independent binned measured SnIa distance moduli appear to be slightly higher than the expected with respect to the best fit Pantheon+ $\Lambda$CDM residual band (the lower stripe in the figure).
  • This tension is less significant for redshift $z>1$ where the SnIa distance moduli residuals with respect to Planck18/$\Lambda$CDM increase while the BAO distance moduli residuals appear to decrease. This reduction of the tension at high $z$ may be viewed as a hint for an $H(z)$ deformation with respect to Planck18/$\Lambda$CDM which however by itself is not sufficient to fully resolve the Hubble tension. This deformation is better expressed by the best fit $\Lambda_{\rm s}$ model than by the best fit $wCDM$ model.
  • Both $H(z)$ deformation models, $wCDM$ and $\Lambda_{\rm s}$ interpolate successfully between the Planck18/$\Lambda$CDM distance modulus to recombination and the Pantheon+ $\Lambda$CDM distance moduli measured by the distance ladder calibrated SnIa at low $z$. However they can not fit simultaneously the BAO and SnIa distance moduli measured at intermediate redshifts since these are inconsistent with each other.
  • The $\Lambda_{\rm s}$ model, due to its high $z$ $H(z)$ transition, appears to provide a better fit to the more constraining SnIa data than the smooth $wCDM$ best fit model. This seems to explain its overall improved quality of fit to the data compared to other $H(z)$ deformation models that attempt to resolve the Hubble tension.
The consistency level between BAO and Pantheon+ SnIa data in the context of $\Lambda$CDM is demonstrated also in Fig. 2. The $\Lambda$CDM $h$-$\Omega_{0m}$ likelihood contours are shown for both BAO and SnIa for two large redshift bins $z\in[0.1,0.8]$ and $z\in[0.8,2.5]$. Clearly the best fit parameter values are in more than $4\sigma$ tension in the low redshift bin while the tension reduces to less than $2\sigma$ in the high redshift bin.
The internal consistency between the high and low redshift bins of the BAO and Pantheon+ SnIa samples is tested in Fig. 3 where we show the parameter likelihood contours for the two bins of the BAO-Pantheon+ samples and for full samples. For the BAO sample the two redshift bins are clearly consistent with each other and with the full BAO dataset while for the Pantheon+ sample there is a mild $2\sigma$ tension between the low redshift bin and the full sample in the direction of $h$. This is also shown in the best fit parameter values with uncertainties shown in Table 2. Finally, the tension between the full BAO and full Pantheon+ SnIa data is more clearly demonstrated in Fig. 4.
IV Conclusion and Discussion
We have explored the level of consistency between the distance moduli of an extensive up to date compilation of BAO and the Pantheon+ SnIa data as a function of redshift. Using a direct comparison of the measured luminosity distance moduli in redshift bins as well as a comparison of the $\Lambda$CDM best fit parameter values in low and high $z$ redshift bins we demonstrated the level of tension between the BAO and SnIa at various redshift bins. This discrepancy is mainly due to the well known inconsistecy of the calibrators used in each sample (recombination sound horizon vs distance ladder calibration of SnIa with Cepheids) known as the Hubble tension. Since we have used data at intermediate redshifts we were able to identify the redshift dependence of this tension as expressed through BAO and SnIa data. In this context our analysis may be viewed as a redshift tomography of the Hubble tension.
We found that the Hubble tension has more prominent presence in relatively lower redshifts $z\in[0.1,0.8]$. This effect is not only due to the higher precision of both the BAO and SnIa data in this redshift range. It is also due to a trend of the best fit values of both distance moduli and $\Lambda$CDM best fit parameters to become more consistent between BAO and SnIa at higher redshift bins ($z\in[0.8,2.5]$. This improved consistency between BAO and SnIa data at higher redshift compared to lower redshifts in the context of distance moduli and $\Lambda$CDM best fit parameters, may also be interpreted as a high $z$ deformation of $H(z)$ with respect to $\Lambda$CDM .
Even though, this deformation can be well expressed by properly designed models like $\Lambda_{\rm s}$ , it is clear from our results that it can not by itself provide good fit to all data. This is due to the calibrator induced inconsistency between the BAO and Pantheon+ SnIa data which is more prominent at the redshift range $z\in[0.1,0.8]$  [^2] This inconsistency between BAO and SnIa distance moduli was first pointed out in Ref. Pogosian et al. (2022) using older BAO and SnIa data, even though no redshift tomography was implemented there..
Interesting extensions of the present analysis include the consideration of additional data like Cosmic Chronometers or multimessenger standard siren data for the determination of the distance moduli in the redshift range considered, This direct comparison with the measured BAO and SnIa measured distance moduli could provide hints about the accuracy of the assumed calibrators (sound horizon vs Cepheid based distance ladder)
Numerical analysis files: The Mathematica v13 notebooks and data that lead to the construction of the figures of the paper may be downloaded from this url.
Acknowledgements
This article is based upon work from COST Action CA21136 - Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), supported by COST (European Cooperation in Science and Technology). This project was also supported by the Hellenic Foundation for Research and Innovation (H.F.R.I.), under the ”First call for H.F.R.I. Research Projects to support Faculty members and Researchers and the procurement of high-cost research equipment Grant” (Project Number: 789).
V Appendix
Here we explain the way in which we obtained the $D_{M}/r_{d}$ measurements of Table 1 for the datapoints with *:
For the measurements 1 and 9, Chuang et al. (2017) provides the following measurements of $D_{A}(z)*(r_{s,fid}/r_{s})$: $D_{A}(0.32)*(r_{s,fid}/r_{s})=956\pm 28Mpc$ and $D_{A}(0.59)*(r_{s,fid}/r_{s})=1421\pm 23Mpc$. Since $D_{M}=D_{A}(z+1)$,to obtain the $D_{M}(z)/r_{d}$ values shown on the table we simply need to multiply by $(r_{s}/r_{s,fid})$ to obtain the measurement of $D_{A}$ and then to multiply by $((z+1)/r_{d})$. The fiducial value $r_{s,fid}$ in Chuang et al. (2017) is $r_{s,fid}=147.66$ and we use $r_{d}=147.18$ and $r_{s}=147.09$. We do the same to obtain the uncertainty of the measurement.
For the measurements 4, 10 and 15, Blake (2012) provides the following measurements of $D_{A}(z)$: $D_{A}(0.44)=1205\pm 114Mpc$, $D_{A}(0.60)=1380\pm 95Mpc$ and $D_{A}(0.73)=1534\pm 107Mpc$. To obtain $D_{M}/r_{d}$ we simply need to multiply by $((z+1)/r_{d})$. We do the same to obtain the uncertainty of the measurements. Seo (2012) (measurement 7) and Delubac et al. (2015) (measurement 30) also provide measurements of $D_{A}$ and thus we follow the same steps. Seo (2012) provides the following: $D_{A}(0.54)=1411\pm 65Mpc$. Delubac et al. (2015) provides the following: $D_{A}(2.34)=1662\pm 96Mpc$.
[…]